OpenStudy (mony01):

By using integration by parts. Solve the solution for the following DE. y'=(t^2-5)cos2t

OpenStudy (tkhunny):

Well, do it? What's stopping you? You \(may\) wish to break it up: \(t^{2}\cos(2t) - 5\cos(2t)\).

OpenStudy (mony01):

whats stopping me is that I need help on it

OpenStudy (loser66):

hahaha!! just take integral both sides

OpenStudy (mony01):

after taking the integral is it \[\frac{ t ^{3} }{ 3 }2\sin(2t)-5tsin(2t)(2)\]

OpenStudy (loser66):

Nope

OpenStudy (mony01):

Can someone show me what I did wrong?

OpenStudy (mony01):

@Loser66

OpenStudy (xapproachesinfinity):

where did you stuck?

OpenStudy (mony01):

in the integral

OpenStudy (xapproachesinfinity):

the whole integral? what is you start point?

OpenStudy (xapproachesinfinity):

integration by part! what did you do?

OpenStudy (xapproachesinfinity):

where did you come with this ?\(\frac{ t ^{3} }{ 3 }2\sin(2t)-5tsin(2t)(2)\)

OpenStudy (xapproachesinfinity):

you are abusing the rules of integration with that anti derivatives

OpenStudy (xapproachesinfinity):

Do you know why?

OpenStudy (mony01):

from the equation tkhunny told me

OpenStudy (xapproachesinfinity):

yes but that's not how you integrate product of two function!!! that rule work with one function standing by it self or a function multiplied with a scalar but here you have an operation that multiplies two functions that's why i said you abused the rules!

OpenStudy (mony01):

oh i needed to use integration by parts

OpenStudy (xapproachesinfinity):

Recall that \(d/dx(f.g)(x)=d/x(f(x))g(x)+f(x)d/dx(g(x))\)

OpenStudy (xapproachesinfinity):

yes! by part. and the question is asking you to do so

OpenStudy (mony01):

im not really good with trig

OpenStudy (xapproachesinfinity):

well then you can't really solve this can you? trig is important for calc!

OpenStudy (xapproachesinfinity):

set you integration by part first?

OpenStudy (xapproachesinfinity):

your*

OpenStudy (xapproachesinfinity):

if it helps remember that \(cos(2x)=cos^2x-sin^2x=2cos^2x-1\)

OpenStudy (mony01):

u=t^2 du=2t dv=cos 2t v=2sin(2t) for the first integral

OpenStudy (mony01):

oh v=2cos^2x-1

OpenStudy (xapproachesinfinity):

why is v=2sin(2t) ?

OpenStudy (xapproachesinfinity):

No it is not i'm just telling you that could be of a help

OpenStudy (mony01):

is it 2cosxsinx?

OpenStudy (xapproachesinfinity):

why? do you know your anti derivative or derivatives of trig functions at least cosine and sine?

OpenStudy (mony01):

yes the antiderivative of cos is sin

OpenStudy (xapproachesinfinity):

what's the problem here v=2sin(2t)

OpenStudy (xapproachesinfinity):

yes that correct! however you are dealing with a composite function here correct? cos(2t)

OpenStudy (xapproachesinfinity):

dv=cos(ax) ====> v=sin(ax)/a

OpenStudy (xapproachesinfinity):

that's Chain rule there yes?

OpenStudy (xapproachesinfinity):

if v=2sint(2t) ===> dv=4cos(2t) which is incorrect, yes?

OpenStudy (xapproachesinfinity):

So what should you do

OpenStudy (mony01):

oh is sin(2t)/2

OpenStudy (xapproachesinfinity):

yes!

OpenStudy (xapproachesinfinity):

Now try working out the first one once you finished go back to the second part

OpenStudy (mony01):

so i have to take the integral for sin(2t)/2du

OpenStudy (xapproachesinfinity):

what is integral by parts ?

OpenStudy (xapproachesinfinity):

int (vdu)=uv-int(udv)

OpenStudy (xapproachesinfinity):

Actually you are looking for int(udv) so int(udv)=uv-int(vdu) in your case

OpenStudy (xapproachesinfinity):

Do you got it? u=t^2 dv=cos(2t) that's what you did, yes?

OpenStudy (mony01):

yes it would be \[\frac{ t ^{2}\sin(2t) }{ 2}-\int\limits \frac{ \sin(2t) }{ 2 }(2t)\]

OpenStudy (xapproachesinfinity):

hmm no! look carefully

OpenStudy (xapproachesinfinity):

oh sorry! that's correct

OpenStudy (xapproachesinfinity):

proceed...

OpenStudy (mony01):

so i take the antiderivative of the integral?

OpenStudy (xapproachesinfinity):

you have to repeat the integral by part! this is a problem that you need to work okay you have to push it so you can solve this!

OpenStudy (xapproachesinfinity):

first cancel 2 you have int(tsin(2t))dt

OpenStudy (xapproachesinfinity):

you could have used u sub, but you need to learn this by part first

OpenStudy (xapproachesinfinity):

Are you working this or not! don't just look at it! paper and pen okay hehehe

OpenStudy (xapproachesinfinity):

by the way did you take calc I, II?

OpenStudy (xapproachesinfinity):

hey are you working on it?